OpenStudy (anonymous):

Find the probability. What is the probability that a card drawn from a deck of 52 cards is not a 10? A. 12/13 B. 9/10 C. 1/13 D. 1/10

OpenStudy (anonymous):

P(not a ten) = (the # of non-tens in the deck)/(total # of cards) = 48/52 = 12/13

OpenStudy (anonymous):

thanks @tcarroll010

OpenStudy (anonymous):

uw!

OpenStudy (raden):

u can use the complement event, P(not 10) = 1 - P(10) = 1 - 4/52 = 48/52 = 12/13

OpenStudy (anonymous):

thank you too @RadEn

OpenStudy (anonymous):

Good luck to you in all of your studies and thx for the recognition! @Girl_Ninja

OpenStudy (anonymous):

;-) @tcarroll010 np